OpenStudy (anonymous):

Please help!!! solve the following system of equations.. x+y=4 x+z=4 y+z=4

OpenStudy (anonymous):

This is a simultaneous equation right??!!??

OpenStudy (anonymous):

i think this ques can be solved by using the concept of inverse of matrix

OpenStudy (anonymous):

Okay :D

OpenStudy (anonymous):

Substitution is the easiest. Eq. 2,3 x+z=4 =4=y+z therefore x+z=y+z Therefore x=y therefore in (1) x+y=4 x+x=4 2x=4 x=2 if (x=2) then 2+Y=4, Y=2 if Y=2 Z=2

OpenStudy (anonymous):

wow thank :-) You wrote it out so easily.